EVA Report – February 5th

Crew 188 EVA #9 Summary Report 05FEB2018
Author’s name: Julia DeMarines, Crew Astronomer, Green Hab Operator

Purpose of EVA: Scouting of Crew 187’s recommended exploration area for chlorophyll sample collection and potential site for artistic Performing Astronautics and Bending Horizons projects using the Insta360 Pro Camera.

Location of EVA: “Despacito Rd” via 1101 off of Cow Dung Road UDM27 Coordinates: 519000 E, 4247500 N (the crew has mapped out several waypoints along potential path for confirmation)

Number of EVA Crew: 3
Participants: Ryan Kobrick (EVA CDR), Sarah Jane Pell, Julia DeMarines EVA Commander: Ryan Kobrick

Road(s) and routes per MDRS map:
1. Take East on Entrance Road (Driving)
2. Turn Right into Cow Dung Road (Driving)
3. Turn right on Crew 187’s “Despacito Rd” at 1101 marker (Driving) 4. Park at the end of the road near fence/gate
6. Resume walking towards West
7. Return by the same route

Mode of Travel: ATV and Pedestrian
Vehicles you will be using (if applicable): Blue ATVs 1, 2, 3, and Red ATV 4 Rovers used: None
Duration: 2 hours 41 minutes
EVA Departure Time: 10:26
EVA Return Time: 13:07


We started up our ATV’s and drove staggered to our first waypoint at the intersection of Cow Dung road and “Despacito Rd” at the 1101 marker. Despacito Rd was much rockier than Cow Dung Rd and we advanced with caution around some high slope and eroded out areas. All four of us made it to our destination safely and parked the ATV’s. Next we walked around and explored the region for suitable spots for Dr. Sarah Jane Pell’s Performing Astronautics and Bending Horizons projects. Sarah Jane found a central spot to set up her Insta360º Pro camera and we recorded about 10 minutes of footage of the crew playing with bubbles at varying distances and heights around the camera. This made an excellent test run for Sarah Jane’s project, which she hopes to begin record tomorrow. Julia DeMarines collected several green samples for her Chlorophyll project.

Once the team was satisfied with their collections, footage, and exploration, we decided to head back home. This was not a difficult decision to make, as it was the hottest day on “Mars” since we arrived, and we were all feeling the heat. Not because we are all (obviously) outstandingly attractive but because the spherical helmet has a way of amplifying the Sun’s rays and we were wearing more than one layer of clothing. On the way back on Cow Dung Rd we stopped at a site of interest that we saw on the way out which looked like another excellent area for Sarah Jane’s Performing Astronautics and Bending Horizons project. We got off of our ATV’s and explored around for 10 minutes finding an even better location for filming. We marked a waypoint on Ryan’s GPS and headed back to the base. Once we had arrived we filled up the gas tanks on the ATV’s while we were waiting for Ryan’s dust project to be ready in the airlock. At 13:07 we entered the airlock, at 13:12 we entered the Hab, and at 13:31 we officially ended EVA #9 after removing our equipment, plugging them in, and cleaning our helmets. We will consider dressing lighter underneath our flight suits on future hot days as it was 20ºC by the time we returned.

EVA Report – February 4th

Crew 188 EVA #8 Summary Report 4FEB2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Purpose of EVA: Collecting soil samples for the micrometeorite investigation and for chlorophyll sample collection on the Hab Ridge Road and Hab View.

Location of EVA: Hab Ridge

UDM27 Coordinates: 5175750E 54250500N

Number of EVA Crew: 3

Participants: Tatsunari Tomiyama, Julia DeMarines, Renee Garifi

EVA Commander: Tatsunari Tomiyama

Rovers used: None

Duration: 2 hours 22minutes

EVA Departure Time: 10:21am

EVA Return Time: 12:43pm


For this EVA, we followed the same foot pathway used on EVA#7. We tried to get to the same locations to collect samples for our micrometeorite investigation and for chlorophyll sampling. The EVA request originally stated we would take 4 hours. However, we planned in the briefing to shorten the length because the path was previously walked. The EVA was successfully completed and we collected samples during this EVA.

The EVA started at 10:21. The weather outside the Hab was nice and windy in this morning. There was no major issue climbing to the Hab Ridge Road. One crew member reported a minor headache at the top and rested for a while because the condition was due to the exercise. After contacting to the HabCOM around the north Hab view, we decided to stay 1 hour to collect samples for research which was a shorter time then we planned in our request. We performed the bubble experiment at both Hab view locations (south and north). On the way back to the Hab, we collected some trash nearby the Hab. Overall, we succeeded in our planned objectives.

EVA Report – February 3rd

EVA #7

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Number of EVA crew: 3

Participants: Sarah Jane Pell, Zac Trolley, and Tatsunari Tomiyama
EVA Commander: Sarah Jane Pell

Purpose of EVA: To get dirty and gather data for the dust study. (Primary Objective) To scout the rim for future EVA missions and observe the area from a high vantage point. (Secondary Objective)

Location of EVA: Hab Ridge, Skyline Rim
UDM27 Coordinates: 5175750E 54250500N

Rovers used: None
Duration: 2 hours 11minutes
EVA Departure Time: 13:30
EVA Return Time: 16:19


There were two objectives for this EVA. Primary objective was to collect data for dust study for NASA and secondary objective was to make more waypoints for future EVA studies. Both objectives were successfully completed.

We had difficulty to find 11:03 road on the way to get Skyline Rim. However, we climbed slopes to get the road while carefully monitoring each other. While climbing, residence-in-artist, Sarah Jane Pell, filmed our activity using bubble.

I am assuming this report will be in journalist report today. After arrived on the Hab Ridge Road, we walked to the South Hab viewpoint and found some animal footprints.

After HABCOM confirmed that we arrived at the objective points and took some photos, we returned to the Hab Ridge Road. We tried to search for the original pathway which is 11:03 and we could found the road.

Then we safely returned.

On the way back to the Hab, we collected some garbage on the past river prints and saw an unknown vehicle as mentioned in the report before.

Dust data collection was performed when we returned to the Hab.

EVA Report – February 3rd

EVA #6

Author’s name: Renee Garifi, Executive Officer

Number of EVA Crew: 3
Participants: Ryan Kobrick, Renee Garifi, Julia De Marines
EVA Commander: Renee Garifi

Location: Lith Canyon and Goblins
UDM27 Coordinates: 519500E, 4256500N

Rovers used: Deimos and Curiosity
Duration: 3 hours, 11 minutes
EVA Departure Time: 10:18
EVA Return Time: 13:29


We drove north on Cow Dung Road to Lith Canyon where we collected hilltop soil samples for the astrobiological micrometeorite investigation and hunted the lower areas for chlorophyll samples to collect. Along the way, we mapped our route using Garmin GPS waypoint drops for archive and analysis of sample collection spots. Both objectives were successfully completed. The final waypoint took the crew to the Goblins site to investigate the area for potential future sampling and video.

We left the airlock at 10:18 and departed for our driving route with Renee and Ryan riding in the Deimos rover and Julia following in Curiosity. At 10:33, we passed Cowboy’s Corner and reached the out-of-range point for the long-range radios. Once in the target area of Lith Canyon, we parked the rovers and proceeded by foot to the Goblins site.

We took photos and videos and collected some scientific samples of rock, soil and small green leaves and rocks for chlorophyll analysis. We also shot some photos and videos for our documentary and enjoyed the breathtaking view from the higher points.

We worked as a team to carefully navigate the rough terrain and slowed our pace when the team was showing signs of fatigue. We opted to skip the Marble Ceremony stop on the way back to the Hab in order to save time and departed Lith Canyon to return to the Hab.

We checked back in with HABCOM at 13:04 when we were back in radio range. We completed the EVA at 13:29.

EVA was a success.

EVA Report – February 2nd

Crew 188 EVA #5 Report 2FEB2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Purpose of EVA: To perform waypoint data collection of North to Galileo road and scout for future EVA sights

Location of EVA: Galileo Road, and Cactus Road

UDM27 Coordinates: 518800 E, 4250600 N and 519200 E, 4252000 N

Number of EVA Crew: 3

Participants: Julia De Marines, *Sarah Jane Pell, and Tatsunari Tomiyama

*Sarah Jane subbed in for Renee due to assessing her fatigue level in the morning and determining she needed to rest and postpone her EVA to tomorrow.

EVA Commander: Tatsunari Tomiyama

Rovers used: Spirit and Curiosity

Duration: 2 hours

EVA Departure Time: 10:50am

EVA Return Time: 12:45pm


We were delayed in our start time for this EVA due to having to re-suit Julia in an alternate EVA suit due to suit #2 having a broken switch. The EVA objective was to collect waypoint data for future EVA sights. This EVA expected and planned to collect waypoints for Galileo Road, and several points on the Cactus Road. Overall, we successfully completed the waypoint data collection on the Cactus Road and Galileo Road.

There were 4 check points on the Cactus Road on the planning. However, we made 6 waypoints on the data collection in this EVA due to the clear visualization. We spent more time than we planned to return to the Hab due to the sample collection being performed at the waypoints.

We made 4 waypoints during this EVA. One is at Galileo Road with the Cow Dung Road as we were trained on driving ATV and rovers. Second waypoint was the intersection with Cactus Road on Galileo Road. This waypoint was relatively difficult to recognize the road visually. We went to the East on cactus Road first and made a waypoint at 11:04. Then, we turned back to the north on the Cactus Road. We stopped and made a waypoint at the entrance of Candor Chasma on Cactus road. We kept driving on Cactus Road to the north. We made a final planned waypoint to the Yellow Zebra Rd. Finally, we visited to the north end of the Cactus Road before reversing our route and returning to the Hab.

EVA Report – February 1st

Crew 188 EVA #4 Summary Report 1FEB2018

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Purpose of EVA: To test human endurance on a pedestrian EVA to Phobos Peak and test the 360 video capture in the field.

Location of EVA: Phobos Peak Rock
UDM27 Coordinates: 519260 E, 4250260 N
Number of EVA Crew: 3
Participants: Julia DeMarines (EVA CDR), Sarah Jane Pell, and Ryan Kobrick EVA Commander: Julia DeMarines Rovers used: none

Duration: 4 hours 14 minutes
EVA Departure Time: 10:44am
EVA Return Time: 14:12pm


The EVA had three primary objectives that were all completed today by the crew:
1 – sample potential “green” samples for Julia’s chlorophyll detection project with NASA Ames
2 – shoot Insta-360 Pro high definition footage for Sarah Jane’s Monash Immersive Visualization Platform project
3 – summit Phobos peak while collecting biometric data for Ryan’s EVA metrics study funded by Florida Space Research Program

The crew faced equipment challenges (see Sol 4 Summary Report) but was able to overcome and persevere to meet the mission goals. The investigators have high hopes for the data collected, but regardless experienced a true Martian adventure.

The crew walked to Phobos Peak with two stops to set up the 360 camera, one stop for sampling, and several stops for interesting formations. We summitted the mountain by scrambling up the West face (the side you can see form the Hab) with several safety stops to combat fogging and shoe repair issues. The international trio waved their 3 flags from the summit for a mini photoshoot from the Hab. On the walk back from the successful mission, the crew stopped for a rare opportunity to toss around a bouncy ball on Mars testing their reflexes, spacesuit glove dexterity and coordination. A radio was dropped near the end of the EVA, but using the GPS track the crew quickly recovered the valuable asset. [end]

EVA Report – January 30th

Crew 188 EVA #3 Summary Report 30 January 2018
EVA Date: 30 Jan 2018
Author: Tatsunari Tomiyama

Purpose of EVA: Ground truthing to establish way points for more accurate future missions, as well as to capture “B roll” of the terrain and us walking for use in video productions. A tertiary goal is to put hours on the rover batteries.

Location of EVA: Lith Canyon
UDM27 Coordinates: 518800 E, 4256500 N
Number of EVA Crew: 3
Crew Members Going on EVA: Ryan Kobrick (Commander), Zac Trolley (Crew Engineer) and Tatsunari Tomiyama (Health & Safety Officer)
EVA Commander: Tatsunari Tomiyama
Rovers used: SPIRIT and CURIOSITY
Rover usage: 2
EVA Duration: 3 hour 27 minutes
EVA Departure Time: 10:24am
EVA Return Time: 13:15pm

Time Check:
09:48 Prep Start + Simulation Start
10:16 Airlock
10:21 EVA Start
10:24 ATV Leave
12:30 Arrived Parking Lot
12:41 Additional near-hab duties, photos and videos
13:01 Airlock
13:15 EVA End

Summary: Our ground truthing reconnaissance was planned to explore the “Goblins” near Lith Canyon. However, we had some difficulty finding our GPS waypoints goals and spent a little more time than planned looking for checkpoints. The planned EVA route was to reach the “Goblins” by way of Cow Dung Road passing Pooh’s Corner and all the way to Lith Canyon Rd 1106. We reached the end of Lith Canyon Rd 1106, took photo and video documentation and began our return to the Hab. Additionally, data was collected for Dr. Kobrick’s EVA for a metrics study on human performance physiology and exploration metrics. We set up checkpoints on this EVA at the following locations: Cowboy Corner, Tank Wash, Intersection to Brahe Hwy 1572, intersection to the Quarry Rd, and the intersection to the Lith Canyon 1106.
On the return trip to the Hab, we collected samples of Gypsum between the intersection to the Brahe Hwy 1572 on the Pooh’s Corner and Tank Wash. We will consult with our crew scientist and mission support team to analyze the sample in the science lab. Once the rovers were parked, we walked around the Hab to get more photos before heading back into the airlock.

EVA Report – January 29th

EVA #2

EVA Date: 29 Jan 2018

Crew Executive Officer Reporting

Purpose of EVA: Spacesuit activity recognition and assessment of in-suit performance
Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)
UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members going on EVA:
Renee Garifi (Executive Officer), Julia De Marines (Crew Astronomer), Zac Trolley (Crew Engineer)

Rovers used:
– Opportunity
– Spirit

Rover time usage: 0.5 hours

Duration of EVA: 1.3 hours

EVA departure time: 15:10
EVA return time: 16:42

Time Checks:
Start Prep: 15:32
Airlock Start: 15:10
Roll out: 15:20
Check-In: 15:32
Comms on Overshoot: 15:40
Scrub due to helmet issue: 15:54
Renee in Airlock: 15:58
Zac and Jules return to get Rover: 15:59
Jules in Airlock: 16:15
Zac in Airlock 16:28
Equipment Cleaned and put away: 4:42

Summary: The Crew attempted to locate the Marble Ceremony landmark near the Pooh’s Corner rock area. They attempted to utilize one of the hab GPS units to identify the planned EVA coordinates. We took with us one hab GPS unit, as well as one of the small, laminated MDRS, printed maps to find our waypoints. Our objective was to perform a training EVA to familiarize our crew (in two team groups of three) with EVA operations first-hand from a participant perspective. All members brought cameras for photos and video footage to send to Mission Support. All three rovers used were fully functional and tested in the rough paths of the Martian Terrain on Cow Dung Road.

We experienced an off-nominal situation with one crewmember during this second EVA of the day. All three of the crew experienced a higher-than-normal amount of helmet fogging while walking. It was realized shortly after parking the rovers on cow dung road that we had overshot the EVA destination and were too far north. While walking back to the rovers, one crewmember experienced helmet fog to the point of having impaired visibility and was unable to drive their rover. After waiting a few minutes for the fogging to clear, the crewmember began to have sweat drip down into the eyes and collect in the glasses. They requested assistance from another crewmember to turn the rover around for them. In the process, the crewmember having the fogging issue experienced a total visual loss when their hat fell from their head and into the front of their helmet, completely blocking their view.

When this hat anomaly occurred, the crewmember made the call to be driven back to the hab to receive assistance with the helmet from the HabCom team. All three crewmembers returned on two rovers, dropping the impaired crewmember at the airlock and returning to retrieve the third rover.

The remaining two crewmembers received permission to resume the EVA and made it out to Pooh’s corner safely and completed the marble ceremony. Despite having to return one crewmember to the hab early, this EVA was highly successful for three reasons. First, it provided a non-life threatening emergency that required immediate assistance for the crew to respond to.

There is a long list of reasons why an EVA will terminate early, so we are fortunate that this instance was minor and could be resolved by removing the helmet back at the hab. Secondly, the crew of three proved they could work together in an off-nominal situation to problem solve and quickly make a decision that places the safety of the crew above the objective of the EVA. Lastly, it allowed us to demonstrate our contingency plan for returning to the hab if one of the rovers breaks down since one rover had to remain parked while all crew returned to the hab.

The crew engineer extended their EVA in order to service the ATVs and Rovers that we used during the day.

EVA Report – January 29th

EVA #1

EVA Date: 29Jan2018

Health & Safety Officer Report

Purpose of EVA: Development of Situational Awareness for Spacesuit Activity

Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)

UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members Going on EVA: Ryan Kobrick (Commander), Sarah Jane Pell (Residence-in-Artist), Tatsunari Tomiyama (Health & Safety Officer)

ATV used: 1,2,3

ATV usage: 5 minutes

EVA Duration: 1 hour 7 minutes

EVA Departure Time: 13:30

EVA Return Time: 14:32

Time Check:

13:01 Prep Start + Simulation Start
13:25 Airlock
13:30 EVA Start
13:35 ATV Leave
13:40 Parked
14:02 Arrived Ritual Marbles
14:18 Return to Hub
14:27 Arrived Parking Lot
14:32 Airlock

Summary: Crews attempted EVA to develop situational awareness for space suit performance. We used 3 ATVs to travel to the targeting location; Marble Rituals. When we arrived at the nearest place on Pooh’s corner, we parked ATV and started to walk. We found a fossil and recent animal footprint soon. Once arrived, we take photos and investigated the location. Then, we returned to the Hub. Overall, there was no significant issue to this EVA and we enjoyed the first simulation.

EVA Report – January 24th

EVA Report for 24Jan2018

Author: Oscar Ojeda

Purpose of EVA: Usage of rovers for battery discharge, aerial mapping of terrain, testing of mid-range EVA protocols, testing of the Guache I Rover, Collection of possible halophile samples.

Destination: Intersection of Quarry Road and Cow Dung Road

Coordinates: 518400 E, 4255500 N

Participants: Atila, Danton, David and Oscar

Narrative: One successful EVA was undertaken on this day. A crew of 4 departed the habitat at 11:35, aiming for the two main goals of discharging the rover’s battery and to keep testing the different projects. After taking Cow Dung Rd for about 30 minutes, the crew arrived at Quarry Road, where several aerial shots were taken, testing the 3D mapping software, as well as a first field run of the Guache I rover. On the way back, the crew stopped on two more locations to perform the same activities, taking aerial footage, and testing the Guache I rover performance over different types of terrain. The expedition never left the main road for more than 20 meters. No halophile samples were taken, because no relevant sites were found during the expedition. The batteries of the rovers were discharged, and in particular, opportunity dropped to 55% during the traverse. The crew entered the airlock at 13:55 and finished the EVA.

No incident or anomaly was sustained during the EVA.